write an equation of the line containing the given point and parallel to the given line (6,8);x+6y=7
Answer by jim_thompson5910(35256)

Anything parallel to x+6y = 7 is of the form x+6y = C where C is any number


Since the line passes through (6,8), we know that x = 6 and y = 8, so plug these values in to find C


x+6y=C

(6)+6(8)=C

6+48=C

54=C

C=54


So the answer is x+6y=54

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim